IIT-JEE 1980Question 1
The probability that an event $A$ happens in one trial of an experiment is $0.4$. Three independent trials of the experiment are performed. The probability that the event $A$ happens at least once is
Probability
Solution
$P=1-(0.6)^3=1-0.216=0.784$.